A man swimming in a river with a speed of 100m\min, can swim continuosly for 5min after he is required to take rest of 1min.A shark chases him with speed of 80m\min.find when will they meet each others if theyare moving opposite to the flow of river with speed 20m\min, also calculate distances travelled by them in 10 min.   

Grade:11

1 Answers

Pooja
4 years ago

The relative speed of man is 100-20=80m/min.
The relative speed of shark is 80-20=60m/min.
Distance travelled by shark in 10 min is
60×10=600m
Distance travelled by man in 10 min is
5 min travel +1min rest + 4 min travel
=5×80 + 0 + 4 × 80
=720 m
The man and shark will never meet as the speed of man is more than that of shark even if he takes rest.